OpenStudy (anonymous):

Solve the equation: x^4+x^2-6=0

OpenStudy (anonymous):

How are we supposed to solve it? like for ex: (2x +1) (x+5) btw thats not the answer.

OpenStudy (anonymous):

Try using the Quadratic Formula?

OpenStudy (anonymous):

(x^2-2)(x^2+3)=0 can you finish it?

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

subsitution method: x^4+x^2-6=0 lets say x^2=a, then a^2+a-6=0 x1=-3 x2= 2 x^2=-3 and x^2=2 x= +/- sqrt2

OpenStudy (anonymous):

x1=-3 x2= 2those are not x1 and x2. They are a1 and a2 my bad

OpenStudy (anonymous):

by factoring x^2*x^2=x^4 x^2*3=3x^2 x^2*-2=-2x^2 -2*3=-6 X^4+3x^3-2x^2-6 simplifies to x^4+x^2-6

hero (hero):

\[x^4+x^2-6 = 0\]\[x^4 +3x^2-2x^2 - 6 = 0\]\[x^2(x^2+3)-2(x^2+3) = 0 \]\[(x^2+3)(x^2-2) =0 \]

hero (hero):

Factor by Grouping Method

OpenStudy (anonymous):

3x^2-2x^2 wouldn't that be just x^2?????

hero (hero):

Yes, I substituted x^2 for 3x^2 - 2x^2 Makes it easier to factor

hero (hero):

Let me guess... You still don't get it..

OpenStudy (anonymous):

Ok, thx